# The Indian Pulsar Timing Array Data Release 2: II. Customised Single-Pulsar Noise Analysis and Noise Budget

### v1

@@ 3.3.2 Optimal Fourier harmonics selection
In order to ascertain the efficacy of this methodology, we also performed a brute-force grid-based approach similar to Srivastava2023. However, instead of doing it sequentially for different red noises, we create a $N^{p}$-dimensional grid, where $p\in[1,3]$ represents the number of red noises in the pulsar as per the optimal model. The density of this grid in each dimension is set the same as in Srivastava2023. The posteriors are sampled at each grid point (which corresponds to a particular red noise model with that value of $N_{\text{harm}}$) and evidences are computed using the DYNESTY package. Once all the evidences are computed, we create a table of Bayes Factors to assess the statistical significance of the choice of Fourier bins. Since this framework is highly computationally expensive, we only performed it for a few pulsars to serve as a sanity check for the results of the dropout power-law model. It should be noted that this procedure has an inherent degeneracy between different noise processes, for instance, whether to select $M$ modes for ARN process or DMN process if they have the same Bayes Factors. Srivastava2023 addresses this issue by pointing out that the presence of simultaneous uGMRT observations in Band 3 and Band 5 makes the InPTA data more sensitive to DM noise. While this treatment was justified for the short time baseline of $\sim$$3.5\,\text{yr}$ in InPTA DR1 data (Tarafdar+2022), the same cannot be extended to the present dataset with a time baseline of $\sim$$7.2\,\text{yr}$ (Rana+2025).
It is worth pointing out that while the dropout power-law model significantly reduces the computational cost from $\mathcal{O}(N^{p})$ to $\mathcal{O}(p)$ as it essentially requires a single hyper-parameter for each red noise model, it suffers from similar degeneracies as the brute-force method due to potential spectral leakages (Iraci+2024). For some pulsars with more than one red noise process, we encountered that the posteriors of k_dropbin were constrained only for one of the noise processes while the remaining ones were quite broad. This was replicated in the brute-force analysis as well where the Bayes Factors were all inconclusive for different combinations of noise processes. In such a scenario, we retained the choice of $N_{\text{harm}}=N_{\text{max}}$ for the process giving poorly constrained k_dropbin posteriors to stay coherent with model selection. If the posteriors are constrained, we consider the greatest integer to the median value as $N_{\text{harm}}$. The optimal Fourier harmonics for different red noise processes in the 27 InPTA DR2 pulsars are summarized in Table 5.
@@ 3.3.3 Parameter Estimation
Table 3: The A-D statistic and p-value estimates obtained using the Anderson-Darling Gaussianity test for the normalised post-noise residuals of the 27 InPTA DR2 pulsars. The pulsars highlighted in red show significant non-Gaussianity in the post-noise residuals, whereas the ones marked in boldface show slight deviations (nearly Gaussian), which might be due to the large number of data points in the sample (Gomez+2021), and are less than $2.5$, thus within 95% confidence intervals (Stephens1974; Goncharov+2021). PSR J1125++7819 is a unique case for which the timing residuals themselves, are Gaussian, therefore we have only used $\texttt{EFAC}=1.0$ for this pulsar and the Gaussianity level corresponds to that of the timing residuals only.
@@ 4 Results
The customised SPNA methodology detailed in Section 3 was applied on the sample of 27 MSPs whose precise timing analysis is done by the InPTA consortium and publicly released as part of the second InPTA Data Release (Rana+2025). We use the par and tim files [^8] https://github.com/inpta/InPTA.DR2.git generated after the InPTA-DR2 timing analysis (refer Section 2 for details) after removing all DMX parameters while retaining the remaining timing-model parameters. We re-analyzed the ToAs by fitting for DM and its first and second derivatives (a Taylor-series representation of secular DM variations) while fixing the $n_{\text{earth}}$ at a constant value of zero. The generated par files are then used as starting points in setting up the deterministic timing model (see Section 3.1). We start by performing model selection for all the pulsars and follow the procedure as detailed in Section 3.3 to arrive at the final results.
@@
Out of the 27 pulsars in the sample, 12 pulsars have $T_{\text{span}}\lesssim 2.5\,\text{yr}$ (see Table 5), all of which were not a part of the first data release (Tarafdar+2022; Srivastava2023) as they were newly added to the sample of InPTA pulsars (Rana+2025). Due to this short time baseline, none of them prefer ARN model or even chromatic noise, except for PSR J1944++0907 where the DMN model is present in the optimal model (see Table 2). Although PSR J2302++4442 has an apparent time baseline of $6.67\,\text{yr}$, the substantial data gap effectively places it within this category (see figure 2 in Rana+2025). PSR J1125++7819 shows no preference to any noise model, even the simplest one with a single EFAC, and the original timing residuals were already Gaussian. Therefore, we consider EFAC=1.0 and no SWdet for this pulsar. All the short time baseline pulsars have fully Gaussian  [^9] The null hypothesis in the Anderson-Darling test is that the distribution is Normal. Therefore, a p-value of $\gtrsim 0.05$ shows strong support towards the distribution being Gaussian. post-noise residuals (see Table 3) which shows that they do not really need any red noise model except for the lone case of PSR J1944++0907.
@@
In contrast, the remaining 15 long time baseline pulsars (with $T_{\text{span}}\gtrsim 2.5\,\text{yr}$) show strong preference to red noises but no generic trend in terms of the optimal model preference. Out of these, only 3 pulsars, namely, PSRs J0437-4715, J1909-3744 and J1939++2134 prefer the most complicated model with all forms of red noises, however, only J1909-3744 shows preference for $\dot{n}_{\text{earth}}$ in the solar-wind model. These 12 pulsars also show a wide range of Gaussianity levels in the post-noise residuals. From Table 3, it is evident that 4 of the 15 pulsars (highlighted in red) show significant non-Gaussianity with the A-D statistic beyond 99% confidence interval of $\sim$$3.9$ which is beyond the acceptable limit of $3\sigma$ level Gaussianity (Stephens1974; Goncharov+2021). 8 of the remaining pulsars (highlighted in boldface) show nearly Gaussian post-noise residuals with the A-D statistic less than $2.5$ for all of them, indicating a non-Gaussianity level below $2\sigma$ (Goncharov+2021). We suspect that this may not be an actual statistical evidence towards non-Gaussianity in the post-noise residuals, rather the high sensitivity of the test towards outliers (which essentially affect the tails of the distributions where these tests are sensitive) and the possible dependence of p-value threshold on the size of the dataset (Gomez+2021). Therefore, we regard this subset of pulsars as nearly Gaussian instead of non-Gaussian in our sample. However, 3 pulsars, namely PSRs J1600-3053, J1730-2304 and J1909-3744, are fully Gaussian. The Gaussianity achieved for PSR J1909-3744 in the presence of all white and red noise model components, shows the efficacy of noise modeling in this work.
@@
The consolidated landscape of the InPTA DR2 noise budget for ARN, DMN and FCN processes is shown in Figures 2, 3 and 4, respectively, showcasing the rich diversity of noise processes present in the InPTA DR2 dataset. Table 5 shows the red noise hyper-parameter estimates obtained using the parameter estimation (see §3.3.3) procedure. The distribution of normalised post-noise residuals for all the pulsars are presented in the B.
@@ 4.1 Examination of non-Gaussian pulsars
As discussed in the previous section, 4 pulsars in the sample of 27 show significant non-Gaussianity (beyond 99% confidence interval) in the post-noise residuals. While this is undesirable, it begs for a deeper examination of the possible causes that might lead to such results. In what follows, we speculate the potential factors that could manifest in the form of high levels of non-Gaussianity in the post-noise residuals for each of these pulsars.
In a previous study, Goncharov+2021 reported significant non-Gaussianity for PSRs J0437-4715, J1022++1001 and J1939++2134. Though we get Gaussianity within $2\sigma$ level for the latter two pulsars, we struggle with similar results for PSR J0437-4715. This MSP is one of the most precisely timed pulsars with a distance of about $157\,\text{parsec}$ (Reardon+2024), which is why it is also one of the brightest pulsars in the sample. Due to very high S/N, the single-pulse level variability gets manifested in the measured ToAs leading to strong jitter and hence scatter in the ToAs across the pulse phase (Liu+2012; pbs+21; Kikunaga+2024). Recently, Kikunaga+2024 reported anomalous frequency dependent trend in the chromatic behavior of jitter in this pulsar, potentially pointing towards a turnover between $500\,\text{MHz}$ and $1260\,\text{MHz}$. Moreover, the treatment of EQUAD and ECORR parameters in this work does not explicitly incorporate the frequency dependence and correlation across adjacent channels of pulse jitter, rather accounts for it in an approximate sense. The complexity of jitter in this pulsar cannot be reliably captured by a generic treatment of white noise, as was done in this work, which may ultimately lead to residual non-Gaussianity in the post-noise residuals.
PSR J1012++5307 shows the highest level of non-Gaussianity in the sample. Srivastava2023 reported only the DMN process for this pulsar (see Section 6 for details) which is not preferred in this analysis. We did include the less preferred ARN+DMN model but found similar non-Gaussianity in the post-noise residuals. We are currently investigating potential causes that may lead to this behavior in this pulsar.
@@
PSR J2145-0750 has been recently associated with a potential mode-change event as well as an epoch-wise variability of the ratio between the two peaks in its average profile (Chowdhury et al., under review). Additionally, the presence of pulse jitter (see discussion in Goncharov+2021) and solar-wind effects (Tiburzi+2019; Susarla2024) have also been reported for this pulsar. An interplay of these potential white and red noise sources may result in unmodeled noise in the timing residuals that shows up as non-Gaussianity in the post-noise residuals.
Finally, spatially correlated red-noise due to GWB is expected to be present in all pulsars, albeit with much smaller amplitude. With increasing ToA precision in future telescopes, this would appear as non-Gaussianity in the timing residuals. While the current precision is much poor for this to explain the above results, some low level contribution from this process to non-Gaussianity on these pulsars is also likely. While a detailed review of each of these potential sources of residual non-Gaussianity in these pulsars is beyond the scope of this work, a more advanced methodology for analysing the noise processes in these pulsars is currently being explored (Dwivedi et al, in preparation).
Figure 5: Consolidated plot of the pre-noise (timing) and post-noise residuals for the 27 InPTA DR2 pulsars, shown in blue and red, respectively. The errors in the case of pre-noise residuals are carried forward from the timing results as per InPTA DR2 (see Section 5 and Figures 7 & 8 in Rana+2025 for details) while the errors in the post-noise residuals are scaled with the white noise parameter estimates as per the optimal model for different pulsars (see Table 2). PSR J1125++7819 is a unique case for which the timing residuals themselves, are Gaussian, therefore we have only used $\texttt{EFAC}=1.0$ for this pulsar.
@@ 5 Analysis of SW-cut datasets
As speculated in Section 4.1, the potential cause of non-Gaussianity in PSRs J1022++1001 and J2145-0750 can be residual solar-wind effects that cannot be accounted for by the deterministic solar-wind model used in this work. In order to gain a better understanding on the same, we analysed these pulsars separately, along with PSRs J1744-1134 and J0751++1807 where some residual non-Gaussianity still persists. We also included PSRs J1730-2304 and J1909-3744, even though we get fully Gaussian post-noise residuals for them, as the strong solar-wind signatures are clearly visible in terms of the annual peaks and a secular trend in their DM time series (see figure 6 in Rana+2025). All of these pulsars lie within the medium-to-high ecliptic latitude regime with PSRs J1022++1001, J1730-2304 and J0751++1807 lying in the low ecliptic latitude regime as per Susarla2024, making them suitable candidates for this analysis (see Table 4).
@@
Since the extent of solar-wind impact and its imprint in the DM time series and timing residuals varies from pulsar-to-pulsar, we adopted a customised approach to decide the cutoff threshold in terms of the solar elongation of the pulsar throughout the year, so as to remove the ToAs near solar-conjunction where the impact of solar-wind is maximum. We adopted an upper limit of $30^{\circ}$ on the solar elongation to make sure that the loss in the number of ToAs remains within 20%. We trimmed the dataset iteratively for each pulsar for every $5^{\circ}$ difference in solar elongation and stopped when the visible solar-wind signatures in their DM time series and ToA residuals were removed. The solar-elongation cut thresholds and the corresponding fractional loss in the number of ToAs are shown in columns 2 and 3, respectively, of Table 4. Figure 6 shows the InPTA-DR2 Band 3+5 DM time series for PSR J1909-3744, before and after the SW-cut is applied. After trimming the datasets, the same noise analysis procedure was repeated for these pulsars where we retained the SWdet model in order to take care of any residual solar-wind effects since the cutoff threshold need not be exhaustive for all pulsars.
The SPNA results on the trimmed datasets are summarised in Table 4 where we see that only PSRs J1022++1001 and J1730-2304 show a deviation in the optimal model compared to the original model (see Table 2). Rest of the pulsars show strong preference to the same model as listed in Table 2 showcasing their optimal nature. In order to quantify the difference in the estimates of various red noises, we have calculated the tension metrics (Raveri+2021) for the red noise amplitude and spectral index between the original results and the ones obtained in the current analysis using the tensiometer [^10] https://github.com/mraveri/tensiometer.git package. We also computed the parameter difference posteriors and tension contours using tensiometer. The tension levels between the two datasets in units of $\sigma$ equivalent to a Normal distribution are shown in column 6 of Table 4. The parameter posteriors, difference posteriors and tension contours between the power-law amplitude and spectral index for all the analysed pulsars are presented in C.

### v2

@@ 3.3.2 Optimal Fourier harmonics selection
In order to ascertain the efficacy of this methodology, we also performed a brute-force grid-based approach similar to Srivastava et al. (2023). However, instead of doing it sequentially for different red noises, we create a $N^{p}$-dimensional grid, where $p\in[1,3]$ represents the number of red noises in the pulsar as per the optimal model. The density of this grid in each dimension is set the same as in Srivastava et al. (2023). The posteriors are sampled at each grid point (which corresponds to a particular red noise model with that value of $N_{\text{harm}}$) and evidences are computed using the DYNESTY package. Once all the evidences are computed, we create a table of Bayes Factors to assess the statistical significance of the choice of Fourier bins. Since this framework is highly computationally expensive, we only performed it for a few pulsars to serve as a sanity check for the results of the dropout power-law model. It should be noted that this procedure has an inherent degeneracy between different noise processes, for instance, whether to select $M$ modes for ARN process or DMN process if they have the same Bayes Factors. Srivastava et al. (2023) addresses this issue by pointing out that the presence of simultaneous uGMRT observations in Band 3 and Band 5 makes the InPTA data more sensitive to DM noise. While this treatment was justified for the short time baseline of $\sim$$3.5\,\text{yr}$ in InPTA DR1 data (Tarafdar et al., 2022), the same cannot be extended to the present dataset with a time baseline of $\sim$$7.2\,\text{yr}$ (Rana et al., 2025a).
It is worth pointing out that while the dropout power-law model significantly reduces the computational cost from $\mathcal{O}(N^{p})$ to $\mathcal{O}(p)$ as it essentially requires a single hyper-parameter for each red noise model, it suffers from similar degeneracies as the brute-force method due to potential spectral leakages (Iraci et al., 2024). For some pulsars with more than one red noise process, we encountered that the posteriors of k_dropbin were constrained only for one of the noise processes while the remaining ones were quite broad. This was replicated in the brute-force analysis as well where the Bayes Factors were all inconclusive for different combinations of noise processes. In such a scenario, we retained the choice of $N_{\text{harm}}=N_{\text{max}}$ for the process giving poorly constrained k_dropbin posteriors to stay coherent with model selection. If the posteriors are constrained, we consider the largest integer less than or equal to the median value as $N_{\text{harm}}$. The optimal Fourier harmonics for different red noise processes in the 27 InPTA DR2 pulsars are summarized in Table 5.
@@ 3.3.3 Parameter Estimation
Table 3: The A-D statistic and p-value estimates obtained using the Anderson-Darling Gaussianity test for the normalized post-noise residuals of the 27 InPTA DR2 pulsars. The pulsars highlighted in red show significant non-Gaussianity in the post-noise residuals, whereas the ones marked in boldface show slight deviations (nearly Gaussian), which might be due to the large number of data points in the sample (Gómez-de-Mariscal et al., 2021), and are less than $2.5$, thus within 95% confidence intervals (Stephens, 1974; Goncharov et al., 2021a). PSR J1125++7819 is a unique case for which the timing residuals themselves, are Gaussian, therefore we have only used $\texttt{EFAC}=1.0$ for this pulsar and the Gaussianity level corresponds to that of the timing residuals only.
@@ 4 Results
The customised SPNA methodology detailed in Section 3 was applied on the sample of 27 MSPs whose precise timing analysis is done by the InPTA consortium and publicly released as part of the second InPTA Data Release (Rana et al., 2025a). We use the par and tim files [^10] https://github.com/inpta/InPTA.DR2.git generated after the InPTA-DR2 timing analysis (refer Section 2 for details) after removing all DMX parameters while retaining the remaining timing-model parameters. We re-analyzed the ToAs by fitting for DM and its first and second derivatives (a Taylor-series representation of secular DM variations) while fixing the $n_{\text{earth}}$ at a constant value of zero. The generated par files are then used as starting points in setting up the deterministic timing model (see Section 3.1). We start by performing model selection for all the pulsars and follow the procedure as detailed in Section 3.3 to arrive at the final results.
@@
Out of the 27 pulsars in the sample, 12 pulsars have $T_{\text{span}}\lesssim 2.5\,\text{yr}$ (see Table 5), all of which were not a part of the first data release (Tarafdar et al., 2022; Srivastava et al., 2023) as they were newly added to the sample of InPTA pulsars (Rana et al., 2025a). Due to this short time baseline, none of them prefer ARN model or even chromatic noise, except for PSR J1944++0907 where the DMN model is present in the optimal model (see Table 2). Although PSR J2302++4442 has an apparent time baseline of $6.67\,\text{yr}$, the substantial data gap effectively places it within this category (see figure 2 in Rana et al. (2025a)). PSR J1125++7819 shows no preference to any noise model, even the simplest one with a single EFAC, and the original timing residuals were already Gaussian. Therefore, we consider EFAC=1.0 and no SWdet for this pulsar. All the short time baseline pulsars have fully Gaussian  [^11] The null hypothesis in the Anderson-Darling test is that the distribution is Normal. Therefore, a p-value of $\gtrsim 0.05$ shows strong support towards the distribution being Gaussian. post-noise residuals (see Table 3), which shows that they do not really need any red noise model except for the lone case of PSR J1944++0907.
@@
In contrast, the remaining 15 long time baseline pulsars (with $T_{\text{span}}\gtrsim 2.5\,\text{yr}$) show strong preference to red noises but no generic trend in terms of the optimal model preference. Out of these, only three pulsars, namely, PSRs J0437-4715, J1909-3744 and J1939++2134 prefer the most complicated model with all forms of red noises, however, only J1909-3744 shows preference for $\dot{n}_{\text{earth}}$ in the solar-wind model. These 12 pulsars also show a wide range of Gaussianity levels in the post-noise residuals. From Table 3, it is evident that four of the 15 pulsars (highlighted in red) show significant non-Gaussianity with the A-D statistic beyond 99% confidence interval of $\sim$$3.9$ which is beyond the acceptable limit of $3\sigma$ level Gaussianity (Stephens, 1974; Goncharov et al., 2021a). 8 of the remaining pulsars (highlighted in boldface) show nearly Gaussian post-noise residuals with the A-D statistic less than $2.5$ for all of them, indicating a non-Gaussianity level below $2\sigma$ (Goncharov et al., 2021a). We suspect that this may not be an actual statistical evidence towards non-Gaussianity in the post-noise residuals, rather the high sensitivity of the test towards outliers (which essentially affect the tails of the distributions where these tests are sensitive) and the possible dependence of p-value threshold on the size of the dataset (Gómez-de-Mariscal et al., 2021). Therefore, we regard this subset of pulsars as nearly Gaussian instead of non-Gaussian in our sample. However, three pulsars, namely PSRs J1600-3053, J1730-2304 and J1909-3744, are fully Gaussian. The Gaussianity achieved for PSR J1909-3744 in the presence of all white and red noise model components, shows the efficacy of noise modeling in this work.
@@
The consolidated landscape of the InPTA DR2 noise budget for ARN, DMN and FCN processes is shown in Figures 2, 3 and 4, respectively, showcasing the rich diversity of noise processes present in the InPTA DR2 dataset. Table 5 shows the median ToA uncertainties and red noise hyper-parameter estimates obtained using the parameter estimation (see §3.3.3) procedure. The distribution of normalized post-noise residuals for all the pulsars are presented in the B.
@@ 4.1 Examination of non-Gaussian pulsars
As discussed in the previous section, four pulsars in the sample of 27 show significant non-Gaussianity (beyond 99% confidence interval) in the post-noise residuals. While this is undesirable, it begs for a deeper examination of the possible causes that might lead to such results. In what follows, we speculate the potential factors that could manifest in the form of high levels of non-Gaussianity in the post-noise residuals for each of these pulsars.
In a previous study, Goncharov et al. (2021a) reported significant non-Gaussianity for PSRs J0437-4715, J1022++1001 and J1939++2134. Though we get Gaussianity within $2\sigma$ level for the latter two pulsars, we struggle with similar results for PSR J0437-4715. This MSP is one of the most precisely timed pulsars because it is one of the brightest pulsars in the sample with a distance of about $157\,\text{parsec}$ (Reardon et al., 2024). Due to very high S/N, the single-pulse level variability gets manifested in the measured ToAs leading to strong jitter and hence scatter in the ToAs across the pulse phase (Osłowski et al., 2011; Liu et al., 2012; Parthasarathy et al., 2021; Kikunaga et al., 2024). Recently, Kikunaga et al. (2024) reported anomalous frequency dependent trend in the chromatic behavior of jitter in this pulsar, potentially pointing towards a turnover between $500\,\text{MHz}$ and $1260\,\text{MHz}$. Moreover, the treatment of EQUAD and ECORR parameters in this work does not explicitly incorporate the frequency dependence and correlation across adjacent channels of pulse jitter, rather accounts for it in an approximate sense. The complexity of jitter in this pulsar cannot be reliably captured by a generic treatment of white noise, as was done in this work, which may ultimately lead to residual non-Gaussianity in the post-noise residuals.
PSR J1012++5307 shows the highest level of non-Gaussianity in the sample. Srivastava et al. (2023) reported only the DMN process for this pulsar (see Section 6 for details) which is not preferred in this analysis. We did include the less preferred ARN+DMN model but found similar non-Gaussianity in the post-noise residuals. We are currently investigating potential causes that may lead to this behavior in this pulsar.
@@
PSR J2145-0750 has been recently associated with a potential mode-change event as well as an epoch-wise variability of the ratio between the two peaks in its average profile (Chowdhury et al., 2026). Additionally, the presence of pulse jitter (see discussion in Goncharov et al. (2021a)) and solar-wind effects (Tiburzi et al., 2019; Susarla et al., 2024) have also been reported for this pulsar. An interplay of these potential white and red noise sources may result in unmodeled noise in the timing residuals that shows up as non-Gaussianity in the post-noise residuals.
Finally, spatially correlated red-noise due to GWB is expected to be present in all pulsars, albeit with much smaller amplitude. With increasing ToA precision in future telescopes, this would appear as non-Gaussianity in the timing residuals. While the current precision is much poor for this to explain the above results, some low level contribution from this process to non-Gaussianity on these pulsars is also likely. While a detailed review of each of these potential sources of residual non-Gaussianity in these pulsars is beyond the scope of this work, a more advanced methodology for analyzing the noise processes in these pulsars is currently being explored (Dwivedi et al, in preparation).
Figure 5: Consolidated plot of the pre-noise (timing) and post-noise residuals for the 27 InPTA DR2 pulsars, shown in blue and red, respectively. The errors in the case of pre-noise residuals are carried forward from the timing results as per InPTA DR2 (see Section 5 and Figures 7 & 8 in Rana et al. (2025a) for details) while the errors in the post-noise residuals are scaled with the white noise parameter estimates as per the optimal model for different pulsars (see Table 2). PSR J1125++7819 is a unique case for which the timing residuals themselves, are Gaussian, therefore we have only used $\texttt{EFAC}=1.0$ for this pulsar.
@@ 5 Analysis of SW-cut datasets
As speculated in Section 4.1, the potential cause of non-Gaussianity in PSRs J1022++1001 and J2145-0750 can be residual solar-wind effects that cannot be accounted for by the deterministic solar-wind model used in this work. In order to gain a better understanding on the same, we analyzed these pulsars separately, along with PSRs J1744-1134 and J0751++1807 where some residual non-Gaussianity still persists. We also included PSRs J1730-2304 and J1909-3744, even though we get fully Gaussian post-noise residuals for them, as the strong solar-wind signatures are clearly visible in terms of the annual peaks and a secular trend in their DM time series (see figure 6 in Rana et al. (2025a)). All of these pulsars lie within the medium-to-high ecliptic latitude regime with PSRs J1022++1001, J1730-2304 and J0751++1807 lying in the low ecliptic latitude regime as per Susarla et al. (2024), making them suitable candidates for this analysis (refer Table 4 for the ecliptic latitudes of these pulsars.)
@@
Since the extent of solar-wind impact and its imprint in the DM time series and timing residuals varies from pulsar-to-pulsar, we adopted a customised approach to decide the cutoff threshold in terms of the solar elongation of the pulsar throughout the year, so as to remove the ToAs near solar-conjunction where the impact of solar-wind is maximum. We adopted an upper limit of $30^{\circ}$ on the solar elongation to make sure that the loss in the number of ToAs remains within 20%. We trimmed the dataset iteratively for each pulsar for every $5^{\circ}$ difference in solar elongation and stopped when the visible solar-wind signatures in their DM time series and ToA residuals were removed. The solar-elongation cut thresholds and the corresponding fractional loss in the number of ToAs are shown in columns 2 and 3, respectively, of Table 4. Figure 6 shows the InPTA-DR2 Band 3+5 DM time series for PSR J1909-3744, before and after the SW-cut is applied. After trimming the datasets, the same noise analysis procedure was repeated for these pulsars, where we retained the SWdet model in order to take care of any residual solar-wind effects since the cutoff threshold need not be exhaustive for all pulsars.
The SPNA results on the trimmed datasets are summarized in Table 4 where we see that only PSRs J1022++1001 and J1730-2304 show a deviation in the optimal model compared to the original model (see Table 2). Rest of the pulsars show strong preference for the same model as listed in Table 2 showcasing their optimal nature. In order to quantify the difference in the estimates of various red noises, we have calculated the tension metrics (Raveri and Doux, 2021) for the red noise amplitude and spectral index between the original results and the ones obtained in the current analysis using the tensiometer [^12] https://github.com/mraveri/tensiometer.git package. We also computed the parameter difference posteriors and tension contours using tensiometer. The tension levels between the two datasets in units of $\sigma$ equivalent to a Normal distribution are shown in column 6 of Table 4. The parameter posteriors, difference posteriors and tension contours between the power-law amplitude and spectral index for all the analyzed pulsars are presented in C.